# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v1

@@ X-B Experiment on PGM Problem
The performance comparison among the four model configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
TABLE II: Reasoning accuracies of models on PGM-neutral.
@@
TABLE III: Generalization Results on PGM.
@@ X-C Experiment on Generative RAVEN and PGM Problems
$\times3$ Grid entire:99.6
@@
$\times3$ Grid entire:99.6
@@
Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning for three categories of RPM problems, making it a suitable validation model for generative RPM tasks. Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
@@
Dataset Sub-task Accuracy (%)
@@
Center 91.0/91.2
$\times2$ Grid 85.2/85.0
@@
L-R 87.9/88.5
U-D 87.9/88.4
O-IC 89.0/89.3
O-IG 83.0/84.2
PGM Neutral 84.5
Table V clearly demonstrates that, without relying on any additional supervisory signals, DIO+WORLD has achieved highly competitive results in open-ended generative RPM problems, which strongly confirms the significant research value of the DIO model. In addition, these results also demonstrate that although mutual information does not directly indicate causal relationships, it remains highly valuable in constructing reasoning capabilities and can withstand the tests conducted by models driven by causal chains.
@@ XI Ablation Study
In this section, this paper conducts a series of ablation experiments. The ablation experiments will be conducted focusing on the key parameter settings.
@@ XI-A Ablation Study on DIO
The ablation study on DIO focuses on the design of the network’s feed-forward structure. Within DIO’s progressive pattern induction module, a permutation and combination strategy ($C_{n}^{r}$) is employed to extract progressive patterns from row and column features. This approach aims to model the causal chain details in RPM problems. In this ablation experiment, we will remove the $C_{n}^{r}$ operation. Instead, we will directly input the total of six row-column features for progressive pattern extraction. By doing so, we aim to obscure the structure in DIO that aligns with the causal chain, thereby validating the significance of causal chain modeling as a methodology for network construction. The obscured DIO is denoted as DIO. The reasoning accuracy achieved by DIO on key sub-tasks of RAVEN, I-RAVEN, and PGM is recorded in Table VI.
TABLE VI: Reasoning Accuracies of DIO.
@@
Table VI indicates that constructing RPM solving models based on causal chains is a critically important condition for achieving high reasoning performance.
@@ XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. This paper demonstrates that the number of constructive incorrect options provided to DIO directly influences the tightness of the variational lower bound on mutual information derived by DIO. Therefore, it is essential to conduct ablation experiments focusing on the number ($\beta$-17) of hypothetical options provided by the Brando network. The reasoning accuracies of DIO+Brando on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $\beta$ values are recorded in Table VII.
TABLE VII: Reasoning Accuracy of DIO+Brando under Different $\beta$ Values.
@@
$\beta$=24 $\beta$=22 $\beta$=20 $\beta$=18 No $\beta$
@@
Table VII indicates that even when only a small number of constructive hypothetical options are introduced into DIO, significant performance improvements can already be achieved for DIO.
@@ XI-C Ablation Study on WORLD
TABLE VIII: Reasoning Accuracy of DIO+WORLD under Different $M$ Values.
@@
The number of components directly affects the complexity of the GMM. As shown in Table VIII, when the number of GMM components is relatively small ($M=32$), meaning that WORLD models a distribution with lower complexity, it will have a negative impact on the performance of DIO in the PGM task. This study posits that the effectiveness of the WORLD method depends on whether the GMM it constructs can effectively represent the inherent complexity of the natural distribution of features encoded by DIO for RPM images. Specifically, when $M$ is set to 32, this GMM with a low number of components cannot adequately capture the complexity of the natural distribution of PGM image features. Forcing the distribution of PGM image features with higher complexity into this simplified GMM will inevitably lead to a performance degradation of DIO.
WORLD introduces an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM theoretically ensures that the WORLD method has a considerable scope of effectiveness. However, the phenomenon where WORLD’s modeling of a GMM with a low number of components causes performance degradation in DIO imposes a crucial prerequisite for the validity of WORLD: the GMM modeled by WORLD must be capable of accommodating the natural complexity of the target problem.
Certainly, in traditional deep learning tasks, underfitting is a prevalent and expected phenomenon that arises when neural networks lack sufficient capacity to handle the complexity of problems. However, the WORLD method makes this phenomenon manageable through its $M$-dependent capacity control. Therefore, particular attention must be paid to the parameter $M$ when applying and adjusting this method.
@@ XII Conclusion
  1. 1.
    Causal Chain Modeling
    • First established a complete causal chain analysis framework of “Image $\rightarrow$ Abstract Attributes $\rightarrow$ Progressive Patterns $\rightarrow$ Pattern Consistency $\rightarrow$ Answer”
    • Provided an explainable causal relationship modeling paradigm for deep learning models.
  2. 2.
    Brando Error Mapping Mechanism
    • Developed a trainable hypothetical incorrect option mapping mechanism to explicitly tighten the variational lower bound of mutual information.
    • Creatively introduced mutual information tightening concepts into abstract reasoning frameworks, thereby opening novel avenues for variational inference methodologies.
  3. 3.
    WORLD Feature Modeling
    • Constructed a Gaussian Mixture Model (GMM) for RPM image feature distributions, enabling infinite and semantically-rich incorrect option sampling
    • Empowering DIO to solve open-ended generative RPM problems.
  4. 4.
    DIEGO Causal Correction Innovation
    • Introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO
    • Established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks
[…]
This study’s multi-dimensional innovations not only achieve technical breakthroughs but, more importantly, provide a new paradigm for causal explainability research in deep learning models, holding significant theoretical value and practical implications for advancing artificial intelligence toward higher-order cognitive intelligence.

### v3

@@ X-B Experiment on PGM Problem
The performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
TABLE II: Reasoning Accuracies on PGM-neutral.
@@
TABLE III: Reasoning Accuracies on PGM Generalization Tasks.
@@ X-C Experiment on Generative RAVEN and PGM Problems
$\times3$ Grid entire:99.8
@@
$\times3$ Grid entire:99.8
@@
Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning, making it a suitable validation model for generative RPM tasks. Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
@@
Dataset Sub-task Accuracy (%) Average (%)
@@
Center 98.0/98.2
91.3/92.6
$\times2$ Grid 82.2/82.0
@@
L-R 96.5/96.9
U-D 95.7/96.9
O-IC 94.9/96.5
O-IG 90.5/96.2
PGM Neutral 80.5 80.5
Table V shows that without extra supervisory signals, DIO+WORLD has attained highly competitive results in open-ended generative RPM problems, strongly confirming the significant research value of the WORLD method. Moreover, these results indicate that though mutual information doesn’t directly reveal causal relationships, it’s still highly valuable for building reasoning capabilities and can withstand the tests by progressive-pattern-driven models.
Notably, the results in Table V still have the potential to improve with further training. In the WORLD method, the ability to tighten mutual information emerges in the early training epochs, whereas the generation capability matures only after the loss term $\ell_{\text{Weighted}}$ stabilizes at a low value, necessitating additional training epochs. To avoid excessive resource consumption, training was halted once the results adequately demonstrated WORLD’s effectiveness.
@@ XI Ablation Study
This section presents a series of ablation experiments focusing on key parameter settings and network architecture.
@@ XI-A Ablation Study on DIO
The ablation study on DIO centers on the design of the network’s feed-forward structure. Within its progressive-pattern induction module, DIO employs a permutation and combination strategy ($C_{n}^{r}$) to extract progressive patterns from row and column features, aiming to model the causal-chain details inherent in RPM problems. For ablation, we remove the ($C_{n}^{r}$) operation and instead feed all six row-column features directly into the progressive-pattern extractor. By masking the structure that aligns with the causal chain, we assess the importance of causal-chain modeling in network design. The resulting accuracies on key sub-tasks of RAVEN, I-RAVEN and PGM are reported in Table VI.
TABLE VI: Reasoning Accuracies of masked DIO.
@@
Table VI shows that building RPM-solving models that explicitly incorporate causal chains is crucial for superior reasoning performance.
@@ XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. $\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO. As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential. Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VII.
TABLE VII: Reasoning Accuracies of DIO+Brando under Different $\beta$ Values.
@@
$\beta$=24(8) $\beta$=22(6) $\beta$=20(4) $\beta$=18(2) Solo DIO
@@
Neutral 98.7 98.4 98.4 98.3 98.1
Table VII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved. The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism. The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
@@ XI-C Ablation Study on WORLD
TABLE VIII: Reasoning Accuracies of DIO+WORLD under Different $M$ Values.
@@
The number of components directly affects the complexity of the GMM. WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability. However, as shown in Table VIII, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task. This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features. If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated. Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
In traditional deep learning tasks, underfitting is a prevalent phenomenon arising when neural networks lack sufficient capacity. However, WORLD makes this phenomenon manageable through its $M$-dependent capacity control. Therefore, particular attention is required to the parameter $M$ when applying this method.
@@ XII Conclusion
XII-1 DIO
We propose that structuring neural network architectures around the causal chains inherent in reasoning problems can lead to enhanced reasoning performance. We also suggest that the learning objectives of end-to-end solving models, including DIO, are essentially variational lower bounds of the mutual information between the context matrix and the correct option. Consequently, both performance and learning efficacy are directly impacted by the looseness of these lower bounds.
XII-2 Brando
We have integrated the concept of mutual information tightening into abstract reasoning frameworks, showing that training with more constructive incorrect options serves as a key mechanism for bounding mutual information. This explores novel pathways for variational inference methodologies. Concurrently, we have developed a constructive hypothetical incorrect option mapping mechanism to analytically constrain the variational lower bound of mutual information.
XII-3 WORLD
We constructed a GMM for the distribution of RPM image features, thereby enabling targeted sampling of infinite and semantically rich options. This approach not only assists the DIO model in tightening mutual information but also equips it with the capability to solve open-ended generative RPM problems.
XII-4 DIEGO
We have introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO. We have established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks.
Combining these four innovations yields a complete technical closed loop of “Architecture Design $\rightarrow$ Training Optimization $\rightarrow$ Data Augmentation $\rightarrow$ Causal Rectification”.